OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television?

OpenStudy (anonymous):

An was _____ the size...? I think u left out a word there/

OpenStudy (anonymous):

3/8 the size

OpenStudy (anonymous):

12 = 3/8x x=32 24=3/8x x=64 The brand name tv is 32 inches by 64 inches

OpenStudy (anonymous):

Omfg thank you so much . .

OpenStudy (anonymous):

ur welcome :)